OpenStudy (anonymous):

Carl can paint a room 3 hours faster than Jennifer can. If they work together, they can complete the job in 2 hours. Using complete sentences, explain each step in figuring out how to determine the time it would take Jennifer to complete this job on her own.

TheSmartOne (thesmartone):

\(\Huge\color{Blue}{Welcome~to~Open~Study!!}\) Lets say Jennifer can do the job in x hours then Carl can do it in x+3 hours \(\Large\color{red}{\frac{1}{x+3}+\frac{1}{x}=\frac{1}{2}}\)

OpenStudy (anonymous):

thank you!